OpenStudy (anonymous):

g(z) = (z^6 + (z^2 - 1)^6)^(-3) find the derivative

OpenStudy (anonymous):

use the chain rule

OpenStudy (anonymous):

let \[g(z) = (z^6 + (z^2 - 1)^6)^{-3}=u\] let \[z^6 + (z^2 - 1)^6 = y\] \[\frac{du}{dx} = \frac{du}{dy}\frac{dy}{dx}\] so find those two derivatives and times them together. this is called the chain rule (as already said above)

OpenStudy (anonymous):

\[u = y^{-3}\]